OpenStudy (jkbo):

solve: x^4 - 18^2+32=0

OpenStudy (raden):

18^2 or 18x^2 ?

OpenStudy (jennychan12):

it would only make sense if it were 18x^2 so... (x^2-16)(x^2-2) = 0 x = +/- 4. sqrt(2)

OpenStudy (jkbo):

18x^2

OpenStudy (jkbo):

@jennychan12 thanks so much but I have one more Question 8x^3 - 27=0

OpenStudy (jennychan12):

8x^3 = 27 x^3 = 27/8 x = \[\sqrt[3]{\frac{ 27 }{ 8 }}\]

OpenStudy (jkbo):

thanks